OpenStudy (anonymous):

What is the slope of the line which passes through (–6, 0) and (4, 0)? undefined –2 0 2

OpenStudy (anonymous):

@Valkarie70

OpenStudy (anonymous):

Use this : \[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (ajprincess):

\[slope~m=\frac{ y_2-y_1}{ x_2-x_1}\] use this formula

OpenStudy (anonymous):

\[slope = \frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

Y=0

OpenStudy (anonymous):

because the two points connected are parrallel across the x axis y=0 makes a horizontal line right through the points parrallel to the x axis, i solved by graphing

OpenStudy (ajprincess):

0/number=0 number/0=undefined

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Here : \(y_1 = y_2 = 0\) \(x_2 = 4\) \(x_1 = -6\)

OpenStudy (anonymous):

You have to just plug these values in the formula for slope...